OpenStudy (anonymous):

how to factor x^2-8x=-16

OpenStudy (anonymous):

it can be solved as x^2 - 8x + 16 = 0 x^2 - 4x - 4x + 16 = 0 x(x-4) - 4(x-4) therefore (x - 4)(x - 4) = (x - 4)^2